# Learn from Global Correlations: Enhancing Evolutionary Algorithm via Spectral GNN

### v3

@@ 4.3 Robustness of GNE for Optimal Solution Deviation
[…]
5 Dicussion
Here, we provide a summary statement to underscore the paper’s central focus:
(1) The principal contribution of this paper is not only the introduction of a graph-evolution framework but also an explicit-modeling approach that aims to revolutionize existing EAs. Moreover, the study delineates clear directions for future EA research in terms of interpretability, effectiveness, and stability.
(2) Compared with existing EAs, Evolution via Spectral Graph Representations benefits from modeling global correlations. The constructed graph is then decomposed to generate spectral components, which are optimized through a frequency-based filter. This control facilitates an automatic exploration–exploitation balance, in which high-frequency modes encode population diversity, whereas low-frequency modes encode population consistency.
(3) We propose a novel paradigm that bridges two fields and unveils the inherent connection between EAs and spectral GNNs, thereby fostering mutual advancement. For EAs, spectral tools provide richer population modeling and a principled basis for operator design. For GNNs, evolutionary operators inspire adaptive, exploration-aware spectral filters, which in turn suggest new network architectures.
6 Conclusion

### v4

@@ Introduction
-10 2.60E+04± $_{\text{$\pm$ 7.94E+03}}$ 3.94E-04± $_{\text{$\pm$ 1.63E-04}}$ 1.20E-05± $_{\text{$\pm$ 5.67E-06}}$ 7.77E-18± $_{\text{$\pm$ 9.82E-18}}$ 2.17E-07± $_{\text{$\pm$ 5.60E-07}}$ 4.22E-19± $_{\text{$\pm$ 1.33E-19}}$
-30 3.51E+04± $_{\text{$\pm$ 1.12E+04}}$ 5.49E-04± $_{\text{$\pm$ 2.49E-04}}$ 1.21E-05± $_{\text{$\pm$ 5.81E-06}}$ 9.00E-18± $_{\text{$\pm$ 1.37E-17}}$ 5.84E-08± $_{\text{$\pm$ 6.79E-08}}$ 4.12E-19± $_{\text{$\pm$ 1.12E-19}}$
-50 5.47E+04± $_{\text{$\pm$ 1.19E+04}}$ 4.01E-04± $_{\text{$\pm$ 1.53E-04}}$ 1.41E-05± $_{\text{$\pm$ 8.43E-06}}$ 5.05E-18± $_{\text{$\pm$ 6.04E-18}}$ 4.55E-07± $_{\text{$\pm$ 1.79E-06}}$ 3.86E-19± $_{\text{$\pm$ 1.17E-19}}$
-70 1.04E+05± $_{\text{$\pm$ 2.16E+04}}$ 3.49E-04± $_{\text{$\pm$ 1.49E-04}}$ 1.63E-05± $_{\text{$\pm$ 8.97E-06}}$ 3.79E-17± $_{\text{$\pm$ 9.31E-17}}$ 2.03E-07± $_{\text{$\pm$ 3.91E-07}}$ 4.39E-19± $_{\text{$\pm$ 1.26E-19}}$
-90 1.71E+05± $_{\text{$\pm$ 2.51E+04}}$ 1.31E-04± $_{\text{$\pm$ 7.58E-05}}$ 1.92E-05± $_{\text{$\pm$ 8.10E-06}}$ 5.65E-17± $_{\text{$\pm$ 1.38E-16}}$ 1.15E-07± $_{\text{$\pm$ 2.22E-07}}$ 4.06E-19± $_{\text{$\pm$ 1.00E-19}}$
+10 2.22E+04± $_{\text{$\pm$ 8.65E+03}}$ 3.97E-04± $_{\text{$\pm$ 2.13E-04}}$ 1.31E-05± $_{\text{$\pm$ 6.11E-06}}$ 9.38E-18± $_{\text{$\pm$ 1.45E-17}}$ 2.16E-07± $_{\text{$\pm$ 5.23E-07}}$ 4.09E-19± $_{\text{$\pm$ 1.01E-19}}$
+30 2.10E+04± $_{\text{$\pm$ 8.80E+03}}$ 4.88E-04± $_{\text{$\pm$ 1.60E-04}}$ 1.13E-05± $_{\text{$\pm$ 4.82E-06}}$ 9.81E-18± $_{\text{$\pm$ 1.94E-17}}$ 2.71E-07± $_{\text{$\pm$ 4.49E-07}}$ 4.30E-19± $_{\text{$\pm$ 9.91E-20}}$
+50 2.60E+04± $_{\text{$\pm$ 1.22E+04}}$ 4.60E-04± $_{\text{$\pm$ 1.49E-04}}$ 1.59E-05± $_{\text{$\pm$ 6.11E-06}}$ 7.66E-18± $_{\text{$\pm$ 9.10E-18}}$ 6.21E-07± $_{\text{$\pm$ 2.51E-06}}$ 4.20E-19± $_{\text{$\pm$ 1.20E-19}}$
+70 4.76E+04± $_{\text{$\pm$ 1.60E+04}}$ 3.36E-04± $_{\text{$\pm$ 1.26E-04}}$ 1.60E-05± $_{\text{$\pm$ 7.53E-06}}$ 1.42E-17± $_{\text{$\pm$ 2.63E-17}}$ 1.09E-07± $_{\text{$\pm$ 1.19E-07}}$ 3.79E-19± $_{\text{$\pm$ 6.79E-20}}$
+90 1.10E+05± $_{\text{$\pm$ 2.67E+04}}$ 1.42E-04± $_{\text{$\pm$ 5.28E-05}}$ 1.99E-05± $_{\text{$\pm$ 7.86E-06}}$ 3.19E-17± $_{\text{$\pm$ 7.22E-17}}$ 8.24E-08± $_{\text{$\pm$ 2.55E-07}}$ 3.79E-19± $_{\text{$\pm$ 1.55E-19}}$
Robustness of GNE in Noisy Function Environments
In real-world scenarios, function evaluations are often affected by measurement errors and uncertainties(Qian et al. 2018; Qian, Yu, and Zhou 2018). In this section, we introduce random noise uniformly distributed in the range [0, 1] to the nine benchmark functions to evaluate the performance of GNE under noisy function conditions. This setup tests the algorithm’s robustness when noise is present in the computation of function values.
To demonstrate the balance between overall performance and computational time in GNE, Figure 4 presents the Friedman ranking(Friedman 1937) and average running times of different algorithms on benchmark functions under both noise-free and noisy conditions. GA and DE, two classic and simple algorithms, exhibit the shortest running times in both conditions but achieve the lowest rankings. CMA-ES and RL-SHADE rank in the middle under both conditions but require the longest and second-longest running times, respectively.SDAES achieves the shortest running time in the noise-free condition and ranks second, slightly behind GNE. However, in the presence of noise, its optimization efficiency significantly deteriorates, dropping to the second-to-last position, even performing worse than the simple algorithm DE. In contrast, GNE achieves the highest ranking under both noise-free and noisy conditions while using a moderate amount of computational time. Additionally, GNE demonstrates strong adaptability to the optimization of functions with added noise, maintaining robust performance across varying conditions.
Robustness of GNE for Optimal Solution Deviation
A critical problem in benchmarking and analysis of evolutionary computation methods is that many benchmark functions have optima located at the center of the feasible set, which can create biases in the evaluation of evolutionary algorithms(Kudela 2022).The center-bias in some methods allows them to easily find optima, rendering comparisons with other algorithms that lack such biases ineffective. To address this issue, we study whether the proposed GNE algorithm exhibits overfitting to the optimal solution by testing its performance under varying deviations. The optimal solution is intentionally shifted by deviations of -10, -30, -50, -70, -90, +10, +30, +50, +70, and +90. This setup is designed to test GNE’s robustness when the optimal solution deviates from its original position.Table 2 presents the performance comparison of GNE with other evolutionary algorithms (GA, DE, CMA-ES, SDAES, and RL-SHADE) under different optimal solution deviations in the Sphere function. This highlights the robustness and adaptability of GNE in handling dynamic changes in the solution landscape, thanks to the translation invariance property of its adjacency matrix, which captures global correlations by modeling relative relationships within the population without relying on the boundaries of the solution space.
Conclusion